Solve (1/81)^t = 243^t-2.
Thanks

Answer by fractalier(2101)

Change everything into powers of 3...so from
(1/81)^t = 243^(t-2)
we get
(3^(-4))^t = (3^5)^(t-2)
3^(-4t) = 3^(5t-10)
and thus
-4t = 5t - 10
-9t = -10
t = 10/9